ID |
Date |
Author |
Subject |
Text |
 |
460
|
Mon Jan 8 19:27:36 2024 |
CB, JM | Leak current tests overnight | Left bias in overnight. Leakage current
decreased and stabilised. See attached. Bias
off as we prepare to mount the fees. |
|
459
|
Sun Jan 7 15:38:57 2024 |
CB, JM | Manifolds mostly assembled + bias overnight check | Assembled most of the connectors on the
new (green) manifolds. Two blanking pulgs
stuck when mounting - cannot remove them |
|
458
|
Fri Jan 5 12:29:10 2024 |
JM, CB | Further bias tests | Biased detectors this morning to 150V,
the leakage current for all detectors appears
stable and is ~3.5 uA. We have also replaced |
|
457
|
Thu Jan 4 13:15:10 2024 |
CB, JM | Tent removed - bias test | Removed tent, octagon, Neg cables.
Disconnected and removed Baking
Pi, TC08, thermocouples. |
|
456
|
Wed Oct 18 16:47:36 2023 |
JG, CB | Power outage | Local power outage at the CRYRING today,
while JG was trying to activate the ion pumps.
Unlcear what caused it. |
|
455
|
Tue Oct 17 08:43:47 2023 |
CB | NEG activation -2 | 09:40 - Found temperature and pressure
stablised, but too high to comfortably activate
NEGs closest to detectors. Detector temperature |
    |
454
|
Mon Oct 16 08:28:02 2023 |
CB | NEG activation -1 | 9:00 - Michael's online vacuum pressure
plot is no longer working - an old image
is being displayed. CARME temperatures around |
|
453
|
Sun Oct 15 09:35:03 2023 |
CB | Starting ramp down | 10:40 - Status as attached. Heaters setpoint/read
at 250 C, 150 C, 225 C. Power stable at 53%,
41%, 63% |
 |
452
|
Tue Oct 3 08:48:36 2023 |
TD | Tuesday 3 October | 09.40 CARME IE514 1.79e-8 mbar, ambient temperature
20.7 deg C
|
  |
451
|
Mon Oct 2 08:41:43 2023 |
TD | Monday 2 October | 09.35 CARME IE514 1.95e-8 mbar, ambient temperature
20.8 deg C
|
6x |
450
|
Sun Oct 1 10:24:50 2023 |
TD | Sunday 1 October | 11.20 CARME IE514 3.53e-9 mbar, ambient temperature
20.6 deg C
|
12x |
449
|
Sat Sep 30 10:58:14 2023 |
TD | Saturday 30 September | 10.50 CARME IE514 7.14e-10 mbar, ambient temperature
20.8 deg C
|
9x |
448
|
Fri Sep 29 08:54:29 2023 |
TD | Friday 29 September |
09.50 CARME IE514 9.0e-11 mbar, ambient temperature
21.3 deg C
|
11x |
447
|
Thu Sep 28 09:19:50 2023 |
TD | Thursday 28 September |
10.20 CARME IE514 8.6e-11 mbar
|
 |
446
|
Wed Sep 27 15:56:41 2023 |
TD | Wednesday 27 September |
16.55 CARME IE514 8.2e-11 mbar
|
  |
445
|
Tue Sep 26 10:13:14 2023 |
TD | ELOTECH PID Temperature Controller R4000 manual ( German and English ) | |
 |
444
|
Tue Sep 26 08:08:42 2023 |
TD | Tuesday 26 September | 09.07 IE514 pressure 6.4e-11 mbar, ambient
temperature +20.3 deg C
|
7x |
443
|
Mon Sep 25 16:33:32 2023 |
TD | Removed tape from heaters TC |
17.33 IE514 pressure 6.4e-11 mbar, ambient
temperature +20.8 deg C
|
  |
442
|
Fri Sep 22 12:35:46 2023 |
JM | Bakeout plots | Attached are plots for the previous bakeout
of CARME which may aid in the bakeout to
come in the following week. |
   |
441
|
Wed Sep 13 15:00:39 2023 |
ML | Vacuum and Heater temperature plots from Grafana | Attached are the plotted curves from vacuum
and heater temperature (actual and set values
from ISOHEAT controllers) during the bakeout |
    |